OpenStudy (anonymous):

For her birthday, Sarah’s parents got her a new phone. The phone’s plan costs $59.99 a month and allows for unlimited calls, unlimited web access, and 100 text messages a month. In addition, each text message over the 100 included with the plan is $0.10 . If Sarah’s parents budget $78.00 a month for her cell phone bill, what is the maximum number of text messages that she can send in one month?

OpenStudy (anonymous):

choices?

OpenStudy (texaschic101):

59.99 gives her 100 text messages. 78.00 - 59.99 = 18.01 text messages cost 0.10 per text 18.01/.10 = 180.1 check.. 59.99 + 180.1(.10) = 78 59.99 + 18.01 = 78 78 = 78 (correct) she can send 100 text messages with the $59.99 plan, plus an additional 180 text messages.....totaling 280 text messages per month and make the budget of $78.